# Offline and Distributional Reinforcement Learning for Radio Resource Management

### v1

@@ Abstract
Reinforcement learning (RL) has proved to have a promising role in future intelligent wireless networks. Online RL has been adopted for radio resource management (RRM), taking over traditional schemes. However, due to its reliance on online interaction with the environment, its role becomes limited in practical, real-world problems where online interaction is not feasible. In addition, traditional RL stands short in front of the uncertainties and risks in real-world stochastic environments. In this manner, we propose an offline and distributional RL scheme for the RRM problem, enabling offline training using a static dataset without any interaction with the environment and considering the sources of uncertainties using the distributions of the return. Simulation results demonstrate that the proposed scheme outperforms conventional resource management models. In addition, it is the only scheme that surpasses online RL and achieves a $16\%$ gain over online RL.
@@ I Introduction
These challenges motivate the importance of offline RL and distributional RL. The former suggests optimizing the optimum policy using a static dataset without any interaction with the environment [8]. In contrast, the latter considers the environment uncertainties relying on the return distribution instead of the average return [9].
@@ II System Model
Consider a wireless network of area $L\times L$ square meters that comprises $N$ randomly deployed access points (APs) and $M$ randomly deployed user equipment (UEs) as shown in Fig. 1. To ensure practical scenarios, the random deployment of APs and UEs are controlled by minimum AP-UE distance $d_{0}$ and minimum AP-AP distance $d_{1}$. Our simulator considers an episodic environment consisting of time slots $t\in\{1,2,\cdots,T\}$, where $T$ is the total number of slots in one episode. At each time slot $t$, each UE moves randomly in the network with a speed $v(t)$. The APs and UEs locations are fixed during each episode.
At the beginning of each episode, user association occurs, where each UE is associated with one of the APs. User-association is performed according to the reference signal received power (RSRP) of each user, where UE $m$ is associated to AP $n$, such that $n={\operatorname{arg}\,\max_{i}}\,\text{RSRP}_{mi}$, where $i\in\{1,2,\cdots,N\}$. The RSRP is affected by the transmit power $p_{t}$ and the channel. The channel between UE $n$ and AP $m$ is characterized by indoor path loss, log-normal shadowing with standard deviation $\sigma_{sh}$, and short-term frequency-flat Rayleigh fading. According to 3GPP [11], the indoor path loss between UE $m$ and AP $n$ is calculated as
@@
$y_{m}(t)=h_{mn}(t)x_{n}(t)+\sum_{i\neq n}h_{mi}(t)x_{i}(t)+n_{m}(t),$ (2)
where $h_{mn}$ is the channel between UE $m$ and AP $n$ and $n_{m}(t)\sim\mathcal{C}\mathcal{N}(0,n_{o}^{2})$ is the additive white Gaussian noise (AWGN) with variance $n_{o}^{2}$. At time $t$, the instantaneous rate (Shannon capacity) of UE $m$ that is associated with AP $n$ is
@@
RRM problems aim to maximize the average data rates across all users. However, this objective is trivial to be solved in a way that each AP always schedule the user with the best SINR. Therefore, fairness across users must be considered in the RRM objective. In this work, we consider joint optimization of the sum-rate and the $5$-percentile rate. The former is calculated as
@@ II-A Problem Definition
where $\eta$ is a step parameter and $\tilde{R}_{m}(0)=R_{m}(0)$. The PF factor is inversely proportional to the long-term rate of the user, reflecting that the higher the PF factor for a user indicates its need to be served. Therefore, the objective in (8) is simplified as
@@ V-B Simulation Parameters and Dataset
We consider an environment of size $50$ m $\times$ $50$ m, $N=3$ APs and $M=15$ UEs. For the online RL, we build a DQN, where the neural network has $2$ hidden layers with $256$ neurons each. Each episode consists of $T=2000$ time steps. We collect the offline dataset using a behavioral policy from an online DQN agent. In other words, we use the last $20\%$ of the transitions of training an online DQN agent. We simulate a single NVIDIA Tesla V100 GPU using Pytorch framework. All the simulation parameters are shown in Table I.
@@
V-C Simulation Results
Fig. 3 demonstrates the convergence of online RL (DQN) as a function of training episodes. We first observe that the random scheme has the worst Rscore, while TDM and full-reuse show close Rscore values. The sub-optimal scheme ITLinQ has the highest Rscore among all the baseline methods. Moreover, DQN reaches convergence (Rscore = $2.6$) after around $150$ episodes. It outperforms all the baseline schemes, including the sub-optimal scheme ITLinQ, by $20\%$.
@@ V-C Simulation Results
In Fig. 4, we report the convergence of the proposed offline and distributional CQR algorithm compared to multiple offline/distributional RL algorithms, namely, CQL, DQN (in an offline manner), and QR-DQN. This figure also illustrates the online scheme (after convergence) and the baseline methods. Although DQN and QR-DQN achieve higher Rscore than TDM and full-reuse, they fail to converge. In contrast, CQL surpasses the sub-optimal ITLinQ; however, it fails to reach the Rscore as the online RL after convergence. Finally, the proposed CQR algorithm is the only one that can reach convergence offline and outperform the online RL.
@@
Fig. 5 shows the performance of the proposed CQR algorithm compared to other baseline offline RL methods during testing after full training. It reports the sum-rate, $5$-percentile rate, and Rscore using an offline dataset contains $20\%$ of the experience of an online DQN (Fig.5a-5c) and using an offline dataset contains $10\%$ of the experience of an online DQN (Fig.5d-5f), respectively. Regardless of the size of the offline dataset, the CQR algorithm outperforms other offline RL schemes. In addition, the size of the dataset slightly affects the performance of CQR as the rates are higher with a larger dataset, which is often better in quality because it comes from the last experiences seen by a DQN agent, usually good experiences. In addition, a significant gap is recorded between the proposed CQR algorithm and other offline schemes with smaller datasets. This highlights that the proposed model requires less data to achieve reasonable rates than other offline RL methods.
@@ VI Conclusions
In this paper, we developed a novel offline and distributional RL algorithm for RRM. First, we formulated the problem as MDP and then introduced the practical limitations of online RL. Afterward, we introduced the proposed model using a combination of CQL and QR-DQN. Simulation results show that the proposed model achieved a higher Rscore than all the baseline schemes. In addition, it is the only scheme to surpass online RL with a $20\%$ gain in terms of the Rscore. Investigating the RRM problem using offline and distributional multi-agent RL is left for future work.
@@ Acknowledgments
This research was supported by the Research Council of Finland (former Academy of Finland) 6G Flagship Programme (Grant Number: 346208).

### v2

@@ Abstract
Reinforcement learning (RL) has proved to have a promising role in future intelligent wireless networks. Online RL has been adopted for radio resource management (RRM), taking over traditional schemes. However, due to its reliance on online interaction with the environment, its role becomes limited in practical, real-world problems where online interaction is not feasible. In addition, traditional RL stands short in front of the uncertainties and risks in real-world stochastic environments. In this manner, we propose an offline and distributional RL scheme for the RRM problem, enabling offline training using a static dataset without any interaction with the environment and considering the sources of uncertainties using the distributions of the return. Simulation results demonstrate that the proposed scheme outperforms conventional resource management models. In addition, it is the only scheme that surpasses online RL with a $10\%$ gain over online RL.
@@ I Introduction
These challenges motivate the importance of offline RL and distributional RL. The former suggests optimizing the optimum policy using a static dataset without any interaction with the environment [8]. This ensures safe optimization and cost minimization. In contrast, the latter considers the environment uncertainties relying on the return distribution instead of the average return [9]. This ensures mitigation sources of risks by considering the worst-case scenarios (tail of the return distribution).
@@ II System Model
Consider a wireless network of area $L\times L$ square meters that comprises $N$ randomly deployed access points (APs) and $M$ randomly deployed user equipment (UEs) as shown in Fig. 1. To ensure practical scenarios, the random deployment of APs and UEs is controlled by minimum AP-UE distance $d_{0}$ and minimum AP-AP distance $d_{1}$. Our simulator considers an episodic environment consisting of time slots $t\in\{1,2,\cdots,T\}$, where $T$ is the total number of slots in one episode. At each time slot $t$, each UE moves randomly in the network with a speed $v(t)$. The APs and UEs locations are fixed during each episode.
At the beginning of each episode, user association occurs, where each UE is associated with one of the APs. User-association is performed according to the reference signal received power (RSRP) of each user, where UE $m$ is associated to AP $n$, such that $n={\operatorname{arg}\,\max_{i}}\,\text{RSRP}_{mi}$, where $i\in\{1,2,\cdots,N\}$. The RSRP is affected by the transmit power $p_{t}$ and the channel. The channel between UE $m$ and AP $n$, $h_{mn}$, is characterized by indoor path loss, log-normal shadowing with standard deviation $\sigma_{sh}$, and short-term frequency-flat Rayleigh fading. According to 3GPP [11], the indoor path loss between UE $m$ and AP $n$ is calculated as
@@
$y_{m}(t)=h_{mn}(t)\>x_{n}(t)+\sum_{i\neq n}h_{mi}(t)\>x_{i}(t)+n_{m}(t),$ (2)
where $n_{m}(t)\sim\mathcal{C}\mathcal{N}(0,n_{o}^{2})$ is the additive white Gaussian noise (AWGN) with variance $n_{o}^{2}$. At time $t$, the instantaneous rate (Shannon capacity) of UE $m$ that is associated with AP $n$ is
@@
RRM problems aim to maximize the average data rates across all users. However, this objective is trivial to be solved in a way that each AP always schedules the user with the best SINR. Therefore, fairness across users must be considered in the RRM objective. In this work, we consider joint optimization of the sum-rate and the $5$-percentile rate. The former is calculated as
@@ II-A Problem Definition
where $\eta$ is a step parameter and $\tilde{R}_{m}(0)=R_{m}(0)$. The PF factor is inversely proportional to the long-term rate of the user, reflecting that the higher the PF factor for a user indicates its need to be served. Therefore, as proved is [5], the objective in (8) is simplified as
@@ V-B Simulation Parameters and Dataset
We consider an environment of size $500$ m $\times$ $500$ m, $N=4$ APs and $M=24$ UEs. For the online RL, we build a DQN, where the neural network has $2$ hidden layers with $256$ neurons each. Each episode consists of $T=2000$ time steps. We collect the offline dataset using a behavioral policy from an online DQN agent. In other words, we use the last $20\%$ of the transitions of training an online DQN agent. We simulate a single NVIDIA Tesla V100 GPU using Pytorch framework. All the simulation parameters are shown in Table I. To ensure scalability, we validate the proposed algorithm compared to the baselines on $100$ unique test episodes.
V-C Simulation Results
@@
Fig. 3 demonstrates the convergence of online RL (DQN) as a function of training episodes. We first observe that the random scheme has the worst Rscore, while TDM and full-reuse show close Rscore values. The sub-optimal scheme ITLinQ has the highest Rscore among all the baseline methods. Moreover, DQN reaches convergence (Rscore = $1.52$) after around $120$ episodes. It outperforms all the baseline schemes, including the sub-optimal scheme ITLinQ, by $20\%$.
@@
In Fig. 4, we report the convergence of the proposed offline and distributional CQR algorithm compared to multiple offline/distributional RL algorithms, namely, CQL, DQN (in an offline manner), and QR-DQN. This figure also illustrates the online scheme (after convergence) and the baseline methods. Although DQN and QR-DQN achieve higher Rscore than the random scheme, they fail to converge. In contrast, CQL surpasses the conventional schemes and the online RL after convergence. However, it fails to reach the convergence of the proposed CQR algorithm, which outperforms the online RL by $20\%$ in terms of the Rscore. This highlights the benefits of optimizing the distributional return over the average return.
@@
Fig. 5 shows the performance of the proposed CQR algorithm compared to other baseline offline RL methods during testing after full training. It reports the sum-rate, $5$-percentile rate, and Rscore using an offline dataset contains $20\%$ of the experience of a converged online DQN (Fig.5a-5c) and using an offline dataset contains $10\%$ of the experience of a converged online DQN (Fig.5d-5f), respectively. Regardless of the size of the offline dataset, the CQR algorithm outperforms other offline RL schemes. In addition, the size of the dataset slightly affects the performance of CQR as the rates are higher with a larger dataset, which is often better in quality because it comes from the last experiences (good experiences) seen by a DQN agent. In the case of using a smaller dataset in training, the proposed CQR algorithm slightly sacrifices the sum-rate to increase the $5$-percentile rate, which achieves higher Rscores than the CQL algorithm, which prefers the sum-rate over the $5$-percentile rate. In addition, a significant gap is recorded between the proposed CQR algorithm and other offline schemes with smaller datasets. This highlights that the proposed model requires less data to achieve reasonable rates than other offline RL methods.
@@ VI Conclusions
In this paper, we developed a novel offline and distributional RL algorithm for RRM. First, we formulated the problem as MDP and then introduced the practical limitations of online RL. Afterward, we introduced the proposed model using a combination of CQL and QR-DQN. Simulation results show that the proposed model achieved a higher Rscore than all the baseline schemes. In addition, it is the only scheme to surpass online RL with a $10\%$ gain in terms of the Rscore. Investigating the RRM problem using offline and distributional multi-agent RL is left for future work. Additionally, the proposed algorithm could be adopted for other 6G applications, such as secure intelligent reflecting surface (IRS) communication and beamformers optimization in mmWave communications.
@@ Acknowledgments
This work was supported by 6G Flagship (Grant Number 369116) funded by the Research Council of Finland.